how can i solve this system by substitution.
2x + 2y = 0
y = 4x + 0

Answer by checkley71(8403)

y=4x+0 now substitute (4x+0) for y in the other equation we get
2x+2(4x+0)=0
2x+8x=0
10x=0
x=0 solution
y=4*0+0
y=0 solution